OpenStudy (anonymous):

find the partial fraction decomp of (5x+3)/(X^2-3x-10)

OpenStudy (anonymous):

Woo partial fractionss

OpenStudy (anonymous):

Alright so first you want to find the (x-a)(x-b) form of the bottom)

OpenStudy (anonymous):

After which you can assume that (5x+3)/(x^2-3x-10) = A/(x-5) + B/(x+2)

OpenStudy (anonymous):

Next you want to multiply everything by the x^2 denominator term to get to 5x+3=A(x+2) + B(x-5)

OpenStudy (anonymous):

Plug in x=-2, x=5 to get your terms for A and B respectively and put them in to the A/(x-5) + B/(x+2) and there's your answer:)

OpenStudy (anonymous):

im lost sry I was out for a month with menijitis

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

so its 7/(x-5) + 2/(x+2) ?

OpenStudy (anonymous):

Going off of where you have things now. When x = -2, you have: 5(-2) + 3 = A(-2+2) + B(-2-5) -7 = -7B B = 1 When x is 5 you have: 5(5) + 3 = A(5+2) + B(5-5) 28 = 7A A = 4 So using those values for A and B, they are plugged into the first fractions that were invented involving A and B, namely the A/(x-5) and B/(x+2), so this means the decomp is 4/(x-5) + 1/(x+2)

OpenStudy (anonymous):

you want a quick way to do this?

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

\[\frac{5x+3}{(x-5)(x+2)}=\frac{A}{x-5}+\frac{B}{x+2}\] if \(x-5=0\) then \(x=5\) put your hand over the factor of \(x-5\) and replace \(x\) by \(5\) in the left hand side \[\frac{5\times 5+3}{\cancel{(x-5)}(5+2)}\] \[=\frac{28}{7}=4\]

OpenStudy (anonymous):

you can do it in your head with a minimum amount of practice repeat with \(x=-2\) \[\frac{5\times (-2)+3}{(-2-5)\cancel{(x+2)}}=\frac{-7}{-7}=1\]

OpenStudy (anonymous):

Cover-up method, hehe.

OpenStudy (anonymous):

thanks a million guys

OpenStudy (anonymous):

just make sure to remember which is A and which is B, it doesn't matter, but you have to keep track also this method only works for simple linear denominator